# Why the Agent Made that Decision: Explaining Deep Reinforcement Learning with Vision Masks

### v1

@@ 3.2 Architecture
As shown in Fig. 1, the expert dataset $\mathcal{D}_{E}$ is collected using the expert policy $\pi_{E}$. From this dataset $\mathcal{D}_{E}$, state-action pairs $(s_{i},a_{i})$ are sampled. Subsequently, the explainer $f_{\theta}$ takes the state $s_{i}$ as input and produces a set of action-wise masks $M=\{m_{0},m_{1},....,m_{K-1}\}$, where $K$ denotes the number of candidate actions and $m_{i}\in{[0,1]^{W\times{H}}}$. A mask value $m_{i}[x,y]$ near 1 indicates that the pixel $(x,y)$ has significant contributions to the corresponding action $i$, whereas a value near 0 denotes irrelavance. This process shares similarities with image segmentation, where each pixel is attributed to an object or background. Hence, we structure the explainer $f_{\theta}$ akin to the widely used image segmentation model, DeepLabV3 (Chen et al. 2017), but retrain it with the expert demonstrations $\pi_{E}$. To make sure that the output masks are bounded to the range $[0,1]$, a sigmoid function is applied at the output of $f_{\theta}$
Following (Stalder et al. 2022), we partition the masks in $M$ into two sets: the active set and the inactive set. Unlike (Stalder et al. 2022), our active set has only one mask, $m_{a}$ where $a$ is the action selected by expert policy $\pi_{E}$ll the other masks, $m_{i}$, where $i\neq{a}$, belong to the inactive set. Each mask $m_{i}$ contains the attributions for the current agent state that explain the corresponding action $i$. The explainer will be trained to produce masks for pixels that contribute the most to the policy’s action selection.
We aggregate the masks in the inactive set into a non-target mask $n$ by taking the maximum value pixel-wise. We also calculate a complement mask $\tilde{m_{a}}=\mathbf{1}-m_{a}$ representing the less important regions for the selected action $a$. Then the masks $m_{a}$ and $\tilde{m_{a}}$ are overlaid onto the original state $s$ to generate two masked states $s_{m}$ and $s_{\tilde{m}}$ using the following overlay function:
$s_{m}=s\odot m_{a}+r\odot(\tilde{m_{a}}),s_{\tilde{m}}=s\odot\tilde{m_{a}}+r\odot(m_{a})$
where $\odot$ is Hadamard Product and $r$ is a reference value. The final step is to feed these masked states $s_{m}$ and $s_{\tilde{m}}$ through the expert policy $\pi_{E}$ to generates action distributions $\mathbf{p}$ and $\mathbf{\tilde{p}}$, respectively:
$\mathbf{p}=\pi_{E}(\mathbf{s}_{\mathbf{m}}),\mathbf{\tilde{p}}=\pi_{E}(\mathbf{s}_{\mathbf{\tilde{m}}}).$
The model $f_{\theta}$ is optimized so that $\mathbf{p}$ matches the action distribution chosen for the unmasked state $s$, while $\mathbf{\tilde{p}}$ does not.
It is worth mentioning that the agent may be sensitive to the reference value in the overlay function. Numerous options exist for the reference value, such as setting the pixel to zero, assigning a constant value, blurring the pixel, or cropping it. In this study, we opt for a constant value as the reference value, based on the sensitivity of the agent to zero and blurring, as indicated by experimental results. We use different reference values for different environments (See Appendix 4).
@@ 3.3 Training Loss
To train the explainer to capture the most important regions in the state, we carefully designed the training loss function $\mathcal{L}$ as follows:
$\mathcal{L}(\mathbf{s},\mathbf{a},\theta,\mathbf{n})$ $=\mathcal{L}_{bc}(\mathbf{s},\mathbf{a})+\lambda_{e}\mathcal{L}_{e}(\mathbf{s},\mathbf{a})+\lambda_{avg}\mathcal{L}_{avg}(\mathbf{s},\mathbf{a},\mathbf{n})$
$+\lambda_{\text{smooth}}\mathcal{L}_{\text{smooth}}(\mathbf{s},\mathbf{a},\mathbf{n})+\lambda_{L2}\lVert\theta\rVert^{2}_{2},$
where $\mathcal{L}_{bc}(\mathbf{s},\mathbf{a})$ is Behavior Cloning loss, $\mathcal{L}_{e}(\mathbf{s},\mathbf{a})$ is entropy loss, $\mathcal{L}_{avg}(\mathbf{s},\mathbf{a},\mathbf{n})$ is the mask average loss, $\mathcal{L}_{\text{smooth}}(\mathbf{s},\mathbf{a},\mathbf{n})$ is the smoothness loss, $\lVert\theta\rVert^{2}_{2}$ is $L_{2}$ regularization and $\lambda_{e},\lambda_{avg},\lambda_{\text{smooth}},\lambda_{L2}$ are regularization hyper-parameters.
Behavioral Cloning loss $\mathcal{L}_{bc}$. Our goal is to train the explainer to select regions in the visual input that that are sufficient and necessary for the agent to choose action $\mathbf{a}$. Hence, we design the behavior cloning loss as cross entropy loss between the action distribution $\mathbf{p}$ predicted for the overlaid state $s_{m}$ and the target $\mathbf{a}$ chosen for the original state $s$:
$\mathcal{L}_{bc}(\mathbf{s},\mathbf{a})=-\frac{1}{K}\sum_{k=1}^{K}\llbracket k=\mathbf{a}\rrbracket\log(\mathbf{p}[k]).$
Here $\llbracket k=\mathbf{a}\rrbracket$ denotes the indicator function which returns 1 if $k$ is the same as label $\mathbf{a}$, and 0 otherwise.
@@
Negative Entropy loss $\mathcal{L}_{e}$. To ensure that the visual input regions selected by $m_{a}$ is sufficient and necessary for the agent to make decisions, we also need to make sure that the unselected region, i.e., $s_{\tilde{m_{a}}}$, does not provide useful information for action selection, hence the action distribution $\mathbf{\tilde{p}}$ should follow a uniform distribution. Motivated by this rationale, we define negative entropy loss regarding $\tilde{\mathbf{m}}$ as the following:
$\mathcal{L}_{e}(\mathbf{s},\mathbf{a})=\frac{1}{K}\sum^{K}_{k=1}\tilde{\mathbf{p}}[k]\log\tilde{\mathbf{p}}[k].$
@@
Mask average loss $\mathcal{L}_{avg}$. A low effort way to minimize $\mathcal{L}_{bc}$ and $\mathcal{L}_{e}$ is to include all pixels in the importance mask, $m_{a}$, and no pixel in the complement mask, $\tilde{m_{a}}$, which obviously is not a valid solution. We need to ensure that each importance mask only consists a small number of crucial pixels. Thus, we define $\mathcal{L}_{avg}$ as follows:
$\mathcal{L}_{avg}(\mathbf{s},\mathbf{a},\mathbf{n})=\frac{1}{W\times H}\sum_{i,j}\mathbf{m}[i,j]+\frac{1}{W\times H}\sum_{i,j}\mathbf{n}[i,j],$
where, $\mathbf{n}$ represents the non-target mask aggregated from the inactive set. By minimizing the mask average loss, we limit the number of pixels selected by the masks.
Smoothness loss $\mathcal{L}_{\text{smooth}}$. Finally, we minimize the Variation between adjacent pixels to generate smoother masks:
$\mathcal{TV}(x)$ $=\frac{1}{W\times H}\sum_{i,j}(|x[i,j]-x[i+1,j]|$
$+|x[i,j]-x[i,j+1]|)$
$\mathcal{L}_{\text{smooth}}(\mathbf{s},\mathbf{a},\mathbf{n})=\mathcal{TV}(m)+\mathcal{TV}(n)$
The overall architecture of VisionMask is given in Fig. 1. The loss will backpropagate through the frozen expert model $\pi_{E}$ to train the explainer model $f_{\theta}$.
@@ 4 Experiments
Env Metric VisionMask LIME RISE Greydanus SARFA
Mario Acc. $\mathbf{0.9076}\pm\scriptstyle{0.0124}$ $0.7868\pm\scriptstyle{0.0027}$ $0.8053\pm\scriptstyle{0.0118}$ 0.1692 0.1692
Precision $\mathbf{0.7109}\pm\scriptstyle{0.0189}$ $0.2448\pm\scriptstyle{0.0200}$ $0.3364\pm\scriptstyle{0.0046}$ 0.6154 0.6154
Recall $\mathbf{0.7005}\pm\scriptstyle{0.0020}$ $0.2467\pm\scriptstyle{0.0042}$ $0.3245\pm\scriptstyle{0.0069}$ 0.1752 0.1800
F1-score $\mathbf{0.6995}\pm\scriptstyle{0.0058}$ $0.2277\pm\scriptstyle{0.0056}$ $0.3046\pm\scriptstyle{0.0082}$ 0.2764 0.2753
Enduro Acc. $\mathbf{0.9061}\pm\scriptstyle{0.0077}$ $0.8811\pm\scriptstyle{0.0087}$ $0.9033\pm\scriptstyle{0.0047}$ 0.2744 0.2790
Precision $0.6058\pm\scriptstyle{0.0084}$ $0.4167\pm\scriptstyle{0.0062}$ $0.5363\pm\scriptstyle{0.0054}$ $\mathbf{0.6449}$ 0.6383
Recall $\mathbf{0.5535}\pm\scriptstyle{0.0049}$ $0.4551\pm\scriptstyle{0.0046}$ $0.5245\pm\scriptstyle{0.0072}$ 0.0839 0.0889
F1-score $\mathbf{0.5060}\pm\scriptstyle{0.0082}$ $0.3809\pm\scriptstyle{0.0018}$ $0.4738\pm\scriptstyle{0.0032}$ 0.1283 0.1418
Seaquest Acc. $\mathbf{0.9531}\pm\scriptstyle{0.0030}$ $0.9094\pm\scriptstyle{0.0079}$ $0.9201\pm\scriptstyle{0.0051}$ 0.4463 0.4476
Precision $0.5809\pm\scriptstyle{0.0022}$ $0.2647\pm\scriptstyle{0.0066}$ $0.4631\pm\scriptstyle{0.004}$ 0.8083 $\mathbf{0.8371}$
Recall $\mathbf{0.4880}\pm\scriptstyle{0.0013}$ $0.2423\pm\scriptstyle{0.0082}$ $0.2830\pm\scriptstyle{0.0047}$ 0.0945 0.0983
F1-score $\mathbf{0.4714}\pm\scriptstyle{0.0050}$ $0.2073\pm\scriptstyle{0.01443}$ $0.2733\pm\scriptstyle{0.0110}$ 0.1640 0.1701
MsPacman Acc. $\mathbf{0.9768}\pm\scriptstyle{0.0177}$ $0.8451\pm\scriptstyle{0.0294}$ $0.9264\pm\scriptstyle{0.0220}$ 0.4616 0.4235
Precision $0.7671\pm\scriptstyle{0.0160}$ $0.1878\pm\scriptstyle{0.0298}$ $0.5590\pm\scriptstyle{0.0097}$ 0.8255 $\mathbf{0.9102}$
Recall $\mathbf{0.8658}\pm\scriptstyle{0.0039}$ $0.1726\pm\scriptstyle{0.0036}$ $0.5587\pm\scriptstyle{0.0035}$ 0.2189 0.1913
F1-score $\mathbf{0.7699}\pm\scriptstyle{0.0096}$ $0.1597\pm\scriptstyle{0.0086}$ $0.5122\pm\scriptstyle{0.0078}$ 0.3307 0.3067
Table 2: Fidelity scores on four environments and confusion matrices in the Mario environment. We run each experiment 3 times with different random seeds and report the mean and standard deviation.
We evaluated VisionMask across four different environments and compared its performance against four baseline methods.
Environment Selection. Four agent-centered RL environments were selected to showcase the effectiveness of VisionMask: Super Mario Bros (SMB)(Kauten 2018), and the Atari games Seaquest, Enduro, and MsPacman. We select these environments following existing works (Puri et al. 2020; Greydanus et al. 2018).
Baseline Selection. We compared VisionMask’s performance against four perturbation-based explainer, LIME (Ribeiro, Singh, and Guestrin 2016) RISE (Petsiuk, Das, and Saenko 2018), Greydanus et al. (2018) and SARFA(Puri et al. 2020). For LIME, we used the original public implementation with the default configuration available from (Ribeiro 2017). For RISE, we adopt the public implementation from torchray (Fong, Patrick, and Vedaldi 2019). For Greydanus et al. (2018) and SARFA, we used the respective public implementation provided by the authors. To ensure a fair comparison, we did not include any explainers that require retraining the agent model.
Expert Policy. For reproducibility, we did not train the expert policies ourselves; instead, we used publicly available implementations. In the SMB environment, We apply the trained PPO (Schulman et al. 2017) policy from (Nguyen 2020). In the Atari environments, we used three trained PPO policies from Stable-Baselines3 (Raffin et al. 2021) .
4.1 Assessment of Explanation Fidelity
We quantify fidelity using four metrics: accuracy, precision, recall, and F1-score. Fidelity measures the agreement between the agent’s action selected for the original visual input and the action selected for the masked visual input. In this context, we define the accuracy as the following:
$\text{Accuracy}=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{i=1}^{|\mathcal{D}_{\text{test}}|}\llbracket\pi_{E}(s_{i})=\pi_{E}(s_{i}^{\prime})\rrbracket,$

### v2

@@ 3.2 Architecture
As shown in Figure 1, we first collect the expert dataset $\mathcal{D}_{E}$ using the expert policy $\pi_{E}$. From this dataset $\mathcal{D}_{E}$, state-action pairs $(s_{i},a_{i})$ are sampled and fed to VisionMas $f_{\theta}$ to generate the set of saliency maps $M$.
Generating the saliency map from given visual input is a dense prediction task that shares similarities with image segmentation, where each pixel is assigned a value to indicate whether it belongs to an object or background. Hence, we structure the explainer $f_{\theta}$ akin to the widely used image segmentation model, DeepLabv3 Chen et al. (2017), however, retrain it using self-supervised contrastive learning. To make sure that the output saliency value are bounded to the range $[0,1]$, a sigmoid function is applied at the output of $f_{\theta}$. For each $m_{i}\in M$, we also calculate a complement map $\tilde{m}_{i}=\mathbf{1}-m_{i}$ highlighting the irrelevant regions for the action $i$. Then the masks $m_{i}$ and $\tilde{m}_{i}$ are overlaid onto the original state $s$ to generate two masked states $s_{i}$ and $\tilde{s}_{i}$ using the following overlay function:
$s_{i}=s\odot m_{i}+r\odot(\tilde{m}_{i}),\tilde{s}_{i}=s\odot\tilde{m}_{i}+r\odot(m_{i})$ (1)
where $\odot$ is Hadamard Product and $r$ is a reference value. Numerous options exist for the reference value, such as setting the pixel to zero, assigning a constant value, blurring the pixel, or cropping it. Empirical study shows that setting the reference to the background gives the best results. More details can be found in the ablation study 4.4. See Appendix F for the background values for different environments.
To generate self-supervised contrastive loss to train the model, we query the agent to obtain the corresponding logits $z_{i}=z(s_{i})$ and $\tilde{z}_{i}=z(\tilde{s}_{i})$, where $z_{i},\tilde{z}_{i}\in\mathbb{R}^{K}$ , and the action probability distributions $p_{i}=\text{softmax}(z_{i})$ and $\tilde{p}_{i}=\text{softmax}(\tilde{z}_{i})$, where $p,p_{i}\in[0,1]^{K}$, $0\leq{i}<K$. By concatenating each $p_{i}$ and $\tilde{p}_{i}$, we have the the action probability distributions of each mask $\mathbf{p},\mathbf{\tilde{p}}\in\mathbb{R}^{K\times K}$,
$\mathbf{p}=[p_{1},p_{2},\dots,p_{k}]^{T}\quad\mathbf{\tilde{p}}=[\tilde{p}_{1},\tilde{p}_{2},\dots,\tilde{p}_{k}]^{T}$
@@ 3.3 Training Loss
To enable the agent contrastively learn the saliency map $m_{a}$, we carefully designed the training loss function $\mathcal{L}$ as follows:
$\mathcal{L}(\mathbf{s},\mathbf{a},\theta)=\mathcal{L}_{a}(\mathbf{s},\mathbf{a})+\lambda_{ne}\mathcal{L}_{ne}(\mathbf{s})+\lambda_{area}\mathcal{L}_{area}(\mathbf{s},\mathbf{a})$
where $\mathcal{L}_{a}(\mathbf{s},\mathbf{a})$ is action-wise contrastive action loss, $\mathcal{L}_{ne}(\mathbf{s},\mathbf{a})$ is feature-wise loss, $\mathcal{L}_{area}(\mathbf{s},\mathbf{a},\mathbf{n})$ is the area size loss and $\lambda_{ne},\lambda_{area}$ are regularization hyper-parameters.
Action-wise contrastive loss $\mathcal{L}_{a}$. Let $a$ denote the target action chosen by the agent. Our primary goal is to learn explanations faithful to the $a$, making $(a,p_{a})$ the only positive pair. Furthermore, the explanation must be discriminative, meaning it should clearly distinguish the target action $a$ from all other possible actions. As a result, every other pair of $(a,p_{i\neq a})$ is treated as a negative pair. Then, we compute the cross-entropy loss between these pairs,
$\mathcal{L}_{pos}(\mathbf{s},\mathbf{a})=-\frac{1}{K}\sum_{k=1}^{K}\llbracket k=\mathbf{a}\rrbracket\log(\frac{\text{exp}(z_{aa})}{\sum_{i=1}^{K}\text{exp}(z_{ai})})$ (2)
$\mathcal{L}_{neg}(\mathbf{s},\mathbf{a})=-\frac{1}{K}\sum_{k=1}^{K}\llbracket k=\mathbf{a}\rrbracket\log(\frac{\text{exp}(z_{aa})}{\sum^{K}_{i=1}\text{exp}(z_{ia})})$ (3)
where $z_{ij}$ is the logit for action $j$ when querying the policy with the masked state $s_{i}$. The contrastive action loss $\mathcal{L}_{a}=\mathcal{L}_{pos}+\mathcal{L}_{neg}$. Here $\llbracket k=\mathbf{a}\rrbracket$ denotes the indicator function which returns 1 if $k$ is the same as label $\mathbf{a}$, and 0 otherwise. Note that we do not compute the loss with $\mathbf{p}_{k,i\neq a}$, as ensuring the faithfulness of the target action $a$ is our primary objective here.
@@
Feature-wise loss $\mathcal{L}_{ne}$. To ensure that the visual input regions selected by $m_{i}$ is necessary for the agent to make decisions, we also need to make sure that the unselected region, i.e., $s_{\tilde{m_{i}}}$, does not provide useful information for action selection, hence the action distribution $\mathbf{\tilde{p}}$ should follow a uniform distribution. Motivated by this rationale, we define negative entropy loss regarding $\tilde{\mathbf{m}}$ as the following:
$\mathcal{L}_{ne}(\mathbf{s})=\frac{1}{K^{2}}\sum_{ij}\tilde{\mathbf{p}}_{ij}\log\tilde{\mathbf{p}}_{ij}.$ (4)
@@
Area size loss $\mathcal{L}_{area}$. A low effort way to minimize $\mathcal{L}_{a}$ and $\mathcal{L}_{ne}$ is to include all pixels in the mask, $m_{i}$, and no pixel in the complement mask, $\tilde{m_{i}}$, which obviously is not a valid solution. We need to ensure that each importance mask only consists a small number of crucial pixels. Thus, we define $\mathcal{L}_{area}$ using L1 norm as follows:
$\mathcal{L}_{area}(\mathbf{s})=\frac{1}{K}\sum_{k}(|\frac{1}{Z}\sum_{i,j}m_{k}[i,j]-a_{max}|)$ (5)
where $Z$ it the number of pixels in state.
SMB Enduro Seaquest
Method Acc. Del. $\downarrow$ Ins. $\uparrow$ LLE $\downarrow$ Sp. $\uparrow$ Acc. Del. $\downarrow$ Ins. $\uparrow$ LLE $\downarrow$ Sp. $\uparrow$ Acc. Del. $\downarrow$ Ins. $\uparrow$ LLE $\downarrow$ Sp. $\uparrow$
LIME 78.7 27.0 30.2 69.4 98.3 88.1 36.3 38.0 47.9 90.2 90.1 14.3 21.4 17.3 97.7
RISE 80.5 18.9 38.0 1.9 2.0 90.3 37.6 39.1 0.01 0.5 92.0 8.1 29.0 0.02 1.4
Greydanus 16.9 56.4 20.2 20.8 63.3 27.4 34.6 33.3 0.25 82.9 44.6 14.4 7.88 0.12 75.1
SARFA 16.9 55.6 21.4 68.7 65.5 27.9 33.3 33.4 0.26 77.0 44.8 7.6 8.7 0.13 75.2
Explainer 92.2 23.6 49.0 38.1 81.2 90.2 34.2 33.1 0.22 81.2 95.3 13.4 30.0 0.9 97.5
VisionMask 95.9 20.4 67.6 38.0 82.3 98.7 32.9 41.2 0.5 80.0 99.6 6.4 34.3 0.9 97.6
MsPacman VizDoom Highway
Method Acc. Del. $\downarrow$ Ins. $\uparrow$ LLE $\downarrow$ Sp. $\uparrow$ Acc. Del. $\downarrow$ Ins. $\uparrow$ LLE $\downarrow$ Sp. $\uparrow$ Acc. Del. $\downarrow$ Ins. $\uparrow$ LLE $\downarrow$ Sp. $\uparrow$
LIME 84.5 32.7 37.7 9.1 96.6 72.1 35.1 11.2 22.1 96.4 77.8 20.0 23.64 52.7 80.5
RISE 92.6 26.2 45.7 0.01 0.59 83.8 14.6 14.5 0.02 0.5 82.5 20.0 22.1 0.05 0.6
Greydanus 46.1 39.1 14.2 0.17 57.4 75.2 20.7 17.2 0.12 62.1 92.7 20.9 24.7 0.93 83.8
SARFA 42.4 18.8 15.6 0.13 66.5 76.1 16.7 17.4 0.55 66.6 97.6 20.4 25.6 0.43 83.0
Explainer 97.6 19.6 59.2 0.22 81.2 84.2 14.4 17.6 0.2 65.5 95.0 20.37 24.4 1.5 83.9
VisionMask 98.7 17.0 62.8 0.2 89.5 87.8 14.2 18.1 3.89 47.8 98.1 19.8 25.8 1.14 84.8
Table 1: Quantitative results on SMB, Enduro, Seaquest, MsPacman, VizDoom and Highway of VisionMask against 5 baselines. Five metrics are compared. The faithfulness is measured by Accuracy(Acc.), Deletion(Del.) and Insertion(Ins.) metrics(%); Robustness is measured by Local Lipschitz Estimate (LLE)(%); And Complexity is measured by Sparseness(Sp.)(%). Blue represents second best results.
Figure 2: Qualitative examples of VisionMask and five baselines across three environments. (a)–(c) show the saliency map overlaid on the input image and counterfactuals where regions are removed based on VisionMask’s map. (a) Human: ”Mario moves left to avoid the Piranha Plant.” VisionMask correctly highlights the Plant; removing it changes the action from ’Left’ to ’Jump right’. (b) Human: ”The agent moves left at constant speed due to the front car.” VisionMask identifies both; removing the car allows acceleration instead of ’Right + accelerate’. (c) Human: ”The agent fires because a shark follows.” VisionMask detects the shark and oxygen bar; removing the shark stops firing. Additional examples in Appendix H.
@@ 4 Experiments
In this section, we begin by outlining the experimental setup. We then present quantitative and qualitative analyses to evaluate our approach. Additionally, we provide counterfactual explanations to demo VisionMask’s faithfulness and sensitivity. Finally, we perform an ablation study to assess the contribution of each component.
4.1 Experimental Setup
Environment Selection. We conduct experiments across three types of environments: Super Mario Bros (SMB) Kauten (2018), Enduro, Seaquest and MsPacman Bellemare et al. (2013); Machado et al. (2018) for 2D game, Highway-env Leurent (2018) for autonomous driving simulation and VizDoom Wydmuch et al. (2018) for 3D game.
Baseline Selection. We mainly compare our model with perturbation-based baselines for black-box RL such as Greydanus Greydanus et al. (2018) and SARFA Puri et al. (2020). In addition, we also compared with three techniques originally designed to explain image classifiers, including a learning-based method, Explainer Stalder et al. (2022), and two perturbation-based methods, LIME Ribeiro et al. (2016) and RISE Petsiuk et al. (2018). Although these methods focus on image classification, their main concept is similar to ours: to attribute a given label (action) to a subset of visual features. We use the public implementation from torchray Fong et al. (2019) for RISE and the original published implementations for other baselines. Among all the baselines, the Explainer is the most similar to VisionMask, as both are learning-based approaches. However, VisionMask employs action-contrastive learning and is trained using different regularization. In the experimental results, we demonstrate that these differences significantly enhances VisionMask’s performance.
Expert Policy. We use the open-source pre-trained PPO Schulman et al. (2017) agents from Nguyen (2020) for SMB environment and from stable-baselines3 Raffin et al. (2021) for Enduro, Seaquest and MsPacman environments. We train the DQN Mnih et al. (2013) agents from scratch for the VizDoom and Highway-env environments. See Appendix A for agents details.
Dataset. We collect state-action pairs for each environment and split the data into 80% for training, 10% for validation, and 10% for testing. All results in this section are reported using the test split. We make this dataset publicly available; see Appendix B for details.
4.2 Quantitative Analysis
Metrics. Since there is no ground truth explanations Adebayo et al. (2020), it is crucial to select appropriate metrics to evaluate the trustworthiness of explanations. Designing robust and consistent metrics for XAI remains an unresolved challenge Hedström et al. (2023). To alleviate the inconsistency between metrics and avoid selection bias, we closely follow the LATEC Klein et al. (2024) benchmark to evaluate performance across three dimensions: Faithfulness, Robustness, and Complexity.